OpenStudy (anonymous):

the diagonal of a rectangle measure 61cm.the width of the rectangle is 11cm find the length of the rectangle.

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

is there anyone to help me

OpenStudy (mathmate):

The diagonal cuts the rectangle into two right-triangles, with the hypotenuse = 61 and one of the (shorter) sides 11. Use Pythagoras theorem for the length L, where L^2+11^2=61^2 Solve for L.

OpenStudy (anonymous):

thank you for your help

OpenStudy (anonymous):

here another question can someone help me please. here the question find the length of side a in the right triangle if b=6 and c=10 can someone help

OpenStudy (mathmate):

Would c=10 be a hypotenuse?

OpenStudy (anonymous):

ok thank

OpenStudy (anonymous):

i need to know what would a be

OpenStudy (mathmate):

If c=10 is a hypotenuse, then a=8 (sqrt(10^2-6^2). If it is not, then we need to find the hypotenuse by sqrt(6^2+10^2)=4sqrt(10)

OpenStudy (anonymous):

i have 8 i was just making sure i was correct thank you

OpenStudy (mathmate):

You're welcome!